Back to the NIKA2 main page

Daily Reports

Dec. 6th, Tuesday

10:30-15:00; Stergios: No observations. Sub-reflector out of order, the team is repairing it at the moment. The problem was identified at the backplane which was replaced. High opacity with low wind. We configured and adopted the new run (i.e., run74) and everything seemed to work in order. Data acquisition was also conducted with bad pixels of 30, 7, 51 for the array 1, 2 and 3, respectively.

15:00-21:00; Juan: Started with NIKA2 observations under average weather conditions. After the configuration of the run 74 we performed a series of DAQs and moved to calibration and project 105-22. During the pointing we observed elevation offsets of >100 arcsec. The problem was detected at the spindle which was stacked at 45 arcsec offset during the installation of the sub-reflector.

21:00-03:00; Quang: We observe the project 105-22, Tau225 was about 0.18, P2P fluctuation was 10 Jy. We did 4 scans of B213-16 (s42 ~ s45), after that we did a POINTING, focusing on 0439+360. We did skydip and POINTING to 0529+075 around 22:10 then focusing & calibration. At last, we did 7 scans of B213-16 (s60, s19~s24).

Dec. 7th, Wednesday

03:00-08:00; Jean-François: Finish up project 105-22 (B213-C6) started by Quang. Over the night, the opacity was "high" between 0.3 and 0.45 but the atmospheric fluctuations were fairly low (< abs(+/-10) Jy/b peak-to-peak at 2mm). Start with focus and beam map on 0316+413 (3C84) with stability ±3Jy/b at 2mm. Project 159-22 (NGC1569). During the last hour, project 169-22 (GJ526North) when other projects were too low in elevation. Stop early at 8:00 local time because of POLAMI.

08:00-10:30; Albrecht: (POLAMI)

10:30-15:00; Maintenance + System check

15:00-21:00; Juan: Observations of the EMIR backup project 124-22 under average weather conditions. At 17:30 UTC we stopped due to snowing and high opacity.

21:00-03:00; Quang: No observations due to high tau.

Dec. 8th, Thursday

03:00-08:00; Jean-François: No observations due to high tau/wind.

08:00-11:00; Albrecht: (POLAMI - no observations).

11:00-15:00; Hélène: No observations due to high tau/wind.

15:00-21:00; Juan: No observations due to high tau/wind.

21:00-03:00; Quang: No observations, strong wind (>20m/s) with rain and snow.

Dec. 9th, Friday

03:00-09:00; Jean-François: No observations. Strong wind (>20m/s) with rain and snow. The anemometer seems frozen showing different values from the actual ones.

09:00-15:00; Matthew: No observations. High wind/tau.

15:00-21:00; Juan: No observations. High wind, snowing.

21:00-03:00; Quang: No observations. High wind, fog, snowing.

Dec. 10th, Saturday

03:00-09:00; Stergios: No observations. It's still snowing/freezing rain while the wind is still above the limit. Antenna parked:
attachment:storm.jpeg attachment:logo.jpg

09:00-15:00; Hélène: No observations. The wind decreased but we still have thick clouds/ high tau.

15:00-21:00; Juan: No observations. After 16:00 UTC the weather improved with lower taus and wind, however the operators checked the sub-reflector for possible ice and it looked completely frozen, due to the ice storm from yesterday night. The ice cannot be removed at this stage so we park the antenna again.
attachment:ice1.jpeg attachment:ice2.jpeg

21:00-03:00; Quang: No observations. See comment above.

Dec. 11th, Sunday

03:00-09:00; Stergios: No observations. Heavy rain and strong wind.

09:00-15:00; Hélène: No observations. Strong wind and dense clouds (tau>1 at all times).

15:00-21:00; Juan: No observations. Lower wind values but with high opacity.

21:00-03:00; Frederic: The weather improved, we had lower tau and wind values so we initiated (around 1:15 UTC) observations of the EMIR backup porjects, starting with 120-22.

Dec. 12th, Monday

03:00-09:00; Frederic/Stergios: We continue EMIR observations under project 162-22. Tau increased to ~0.85.

09:00-15:00; Hélène/Stergios: Continuing with EMIR observations and project 162-22 until 10:50 UTC. We switched to project 135-22 after this time. The opacity is high with pwv values given in the calibrations below 10mm at most times. Stopped at 11:30 UTC due to high opacity. No data from project 135-22 could be acquired. Stopped until the end of the shift.

15:00-21:00; Juan: No observations. Strong wind and dense clouds (tau>1 at all times).

21:00-03:00; Quang: No observations. Strong wind and dense clouds.

Dec. 13th, Tuesday

03:00-08:00; Jean-François: No observations. Strong wind and dense clouds.

09:00-10:30; Stergios: No observations. Strong wind and dense clouds.

DailyReportsNika2Pool061222 (last edited 2022-12-13 08:19:02 by NikaBolometer)